Factor the following equation a) f(x) = (x ^ 2 - 36)/(x ^ 2 + 3x + 2) f(x) = 2x ^ 3 + 3x ^ 2 - 3x - 2


1
Expert's answer
2021-02-23T08:22:21-0500

a) Consider the function "f(x)=\frac{x^2-36}{x^2+3x+2}"


Factor the numerator "x^2-36" using "a^2-b^2=(a+b)(a-b)" as,


"x^2-36=(x+6)(x-6)"


Factorize the denominator as,


"x^2+3x+2=x^2+2x+x+2=(x+2)(x+1)"


Therefore, the function in factored form is "f(x)=\frac{(x+6)(x-6)}{(x+2)(x+1)}"


b) Consider the function "f(x)=2x^3+3x^2-3x-2"


For "x=1,f(1)=2+3-3-2=0" , so "x-1" is a factor of the function.


Expand the function to find the factor "x-1" as,


"f(x)=2x^3-2x^2+5x^2-5x+2x-2"


"=2x^2(x-1)+5x(x-1)+2(x-1)"


"=(x-1)(2x^2+5x+2)"


"=(x-1)(2x^2+4x+x+2)"


"=(x-1)(2x(x+2)+1(x+2))"


"=(x-1)(2x+1)(x+2)"


Therefore, the factored form of the function is "f(x)=(x-1)(2x+1)(x+2)" .